# Towards Bidirectional Human-AI Alignment: A Systematic Review for Clarifications, Framework, and Future Directions

### v2

@@ 5. Paper Coding Analyses and Findings
In this section, our aim is to consolidate key findings derived from our analysis of the framework and the current state of literature we reviewed. We begin by analyzing the overall trends and gaps in the literature (Section 5.1). We then focus on three essential aspects: the relationship between human values and alignment (Section 5.2), the potential interaction modes used to specify human values (Section 5.3), and the gaps between AI model and human study evaluation (Section 5.4).
5.1. Overview of Trends and Gaps in the Literature
@@
Based on our coding of all papers, we computed the number of relevant papers for each dimension in the framework (see Figure 8). We noticed that certain dimensions are over- or under-represented. Most literature specified human values using explicit human feedback, whereas implicit and simulated human feedback were largely under-explored. Additionally, many studies on human cognitive adjustment to AI focus on enabling AI sensemaking and explanations so that humans can better understand, trust, and rely on AI. However, these studies often focused on explaining AI decision-making justification rather than educating people to acquire general skills and competencies required to understand, use, critique, and interact effectively with AI systems (i.e., AI literacy). AI literacy (Long and Magerko, 2020) plays a fundamental role in ensuring people understands and uses AI correctly, but it still needs to be explored compared to other topics like explanation and trustworthiness. Also, we observed a wide range of studies (Wu et al., 2022a; Gebreegziabher et al., 2023; Park et al., 2023; Ma et al., 2024) developing interactive mechanisms and prototypes to empower humans to collaborate with AI in diverse roles of AI. However, most existing studies assumed that AI plays an assistant role, being less capable than humans. This situation might change in the long term. Moreover, the influence of AI advancements on human behavior, social relationships, and societal changes is essential but remains largely unexplored.
5.2. Findings in Human Values for Alignment
@@
We present the key findings for future research which are identified during our adjustment of the “Schwartz Theory of Basic Values” and literature review.
5.2.1. The Relative Priority of Values is as Crucial as Their Existence
@@
5.2.2. Human-desired AI Values and Customization to Personal Values
While further research in AI ethics is crucial in this nuanced area, there are values that, in general, humans expect AI to prioritize (e.g., capability, equity, responsibility) and values that humans do not want AI to integrate in specific scenarios (e.g., seeking power to avoid or harm people (Ngo et al., 2024)). On the other hand, AI models are expected to be tailored to accommodate diverse human value systems (Sorensen et al., 2024), which necessitates a comprehensive understanding of all types of values. Therefore, future research could potentially explore the directions of: (1) developing methods to identify the appropriate sets of human values to align with specific individuals or groups, and (2) understanding and customizing AI to align with users’ values while maintaining its inherent ethical principles.
5.2.3. Expectations and Evaluations of Values May Differ Between Humans and AI Systems
Due to the fundamental differences between humans and AI models, the expectations for evaluating certain values can vary between the two. For instance, while human honesty can hardly be assessed except observable actions, for AI models the expectations might be higher. It could involve reverse-engineering the computational mechanisms and representations learned by neural networks into concepts understandable to humans. This approach, known as mechanistic interpretability, provides a in-depth understanding of how AI models operate (Bereska and Gavves, 2024). Future research is needed to explore evaluating AI values and calibrating human expectations of these values.
5.3. Interaction Techniques for Human Value Specification
Selecting suitable interaction techniques for humans to specify AI values for alignment is crucial. Each research domain offers distinct strengths and insights to address this challenge. From the bidirectional perspective, we investigate the status quo of existing interaction techniques for alignment specification in both human-centered (e.g., HCI) and AI-centered (e.g., NLP/ML) studies. Additionally, we provide a detailed analysis of how these techniques have been utilized in the learning and inference stages, as shown in Figure 9.
5.3.1. Interaction Techniques Across Various Domains
As shown in Figure 9, the coverage of interaction techniques in AI-centered (i.e., NLP/ML Research) and human-centered (i.e., HCI Research) studies often differ (Brath, 2021). Particularly, we see that NLP/ML studies primarily leverage numeric-based, natural language-based interaction techniques for human-AI interaction and alignment. In addition, NLP/ML researchers also leverage implicit feedback to improve the model. For example, some systems took in implicit signals such as contextual information, or discarded options to develop AI systems. In comparison, HCI studies cover more diverse graphical multi-modal interaction signals (e.g., sketches, location information) beyond text and images, which can be largely attributed to the proficient graphical user interface development skills found in HCI researchers and practitioners. This potentially indicates gaps in the human behavioral information that the two fields can extract from the users of the AI models.
5.3.2. Different Interaction Techniques in Separate Stages
Another finding we discerned is the different interaction techniques used in separate stages, especially in the NLP/ML fields. Specifically, the learning stage relies mostly on rating and ranking interaction to deal with the generated datasets of the models. However, users leverage the AI model in different ways during inference stage, as demonstrated in HCI research. This indicates that users might elicit different needs in practice when interacting with AI during the model deployment stage. Therefore, future NLP/ML and HCI researchers might collaborate to keep the human-AI interaction techniques informed (or even consistent) by both domains in order to facilitate close alignment between humans and AI.
5.3.3. Dissimilar Usage of Interaction Data Between Domains
We noticed that the ways in which the two domains utilize data differ. For example, the interaction outputs of NLP research are commonly the dataset to be trained in the model, whereas HCI researchers use usage data to analyze human behavior and feedback. As AI systems continue to advance, new modes of interaction are required to capture more space of human expression, with the goal of deeper and meaningful alignment. This is an important area for NLP/ML researchers to partner with HCI researchers for designing interaction paradigms, including but certainly not limited to affective computing, accessibility, social computing, computer-supported cooperated work, and ubiquitous computing.
5.4. Gaps Between Evaluating AI Systems and Human Study Evaluation
Improving bidirectional alignment between humans and AI can potentially be further enhanced by collaborative efforts and commitments from both the NLP/ML and AI domains, ensuring consistency in evaluation methods and metrics. Nevertheless, we observed notable gaps by comparing the literature analysis regarding “evaluating AI systems” (i.e., Section 4.2.3) and “evaluation in human studies” (i.e., Section 4.4.3).
5.4.1. Different Objectives and Expectations of Evaluations in Humans and AI
Evaluating AI systems primarily focuses on what algorithmic performance is, aiming to determine how well proposed methods perform compared with benchmark or human-in-the-loop evaluation. Conversely, human study evaluations are more concerned with why – understanding the user experience and interactions, often prioritizing both quantitative and qualitative measurements to gain a comprehensive understanding of user experience and interactions with AI systems. Given the distinct objectives in evaluation, future research has the potential to collaborate in bridging these gaps. For instance, informed by the user experience and demands revealed by HCI studies, NLP/ML researchers can incorporate the insights and design human-centered evaluation benchmarks to assess AI systems.
5.4.2. Contrasting Ways of Utilizing User Information
Given the diverse objectives of assessment described above, NLP/ML research primarily focuses on quantifiable measures such as accuracy, precision, and recall. The information received in evaluating AI systems is often accurately measurable, e.g., ratings, rankings, and multiple choice questions. In comparison, by focusing on understanding the user experience and interaction with AI systems, HCI research often involves gathering both quantitative and qualitative data, such as user feedback, perceptions, and behaviors. The collected information from HCI research might cover a wider spectrum of qualitative data related to human values, experience, and needs in practice than what can be captured by quantitative metrics alone. However, incorporating qualitative findings into improving AI systems can be difficult. Therefore, future researchers can investigate methods to convert qualitative human feedback into quantitative data that AI systems can effectively utilize and align with.
5.4.3. Evaluation Approaches and Scale
Prior studies in NLP/ML have predominantly relied on automated or human-in-the-loop evaluation methods to assess AI system performance across extensive datasets. In comparison, HCI studies typically employ qualitative and/or quantitative measurements to gain deeper insights into user experiences and interactions within defined human groups. These divergent aspects highlight diverse gaps between the evaluation methods and scales utilized in the two domains. Closing this gap and adopting consistent evaluation strategies could lead to a more holistic understanding and synergy of human-AI interaction and alignment across both fields.
Overall, it’s important to note that the disparities between HCI and NLP/ML studies extend beyond the above aspects, e.g., research methods and problem formation. Our analyses aim to pave the way for future interdisciplinary research collaboration, fostering bidirectional alignment between humans and AI across diverse domains.
@@ 6. Future Directions
The two directions of human-AI alignment are not independent; instead, they represent an ongoing, mutual process to achieve a dynamic alignment goal over the long term. Future research should approach alignment studies as a bidirectional process to fully explore the breadth of the research space and achieve this essential goal.
@@
Particularly, we discuss research questions and motivations followed by ideas for solution exploration from both directions of “
Align AI to Human” (i.e., domains related to AI algorithm and method development, including Machine Learning, Artificial Intelligence, Data Science, and others) and “
Align Human to AI” (i.e., domains conducting studies related to humans and society, including Human-Computer Interaction, Social Science, Cognitive Science and others) for each challenge below.
6.1. Challenge 1: Specification Game
@@ 6.1.1. Aligning AI to Humans: Integrate fully specified human values into aligning AI
Individuals often possess value systems that encompass multiple values with varying priorities, rather than a single value, to guide their behaviors (Schwartz, 1994, 2012). For example, people may prioritize the value of family and work differently, leading to varied choices in time allocation. Also, these priorities can change dynamically throughout an individual’s life stages. As such, it is unlikely to that we can achieve a universal value system, recognized by all humans, to govern AI alignment. Instead, it is more realistic to select values compatible with specific societies or situations, given the fact that we live in a diverse world (Gabriel, 2020). Therefore, in the absence of a universal value agreement, it is crucial yet challenging for AI designers to investigate how to fully specify the appropriate values and to further integrate these values into AI alignment.
Examples of Potential Future Solutions. Future work can draw from political theory, which examines fair methods for determining the principles that state institutes should follow, as a potential guide for selecting human values for AI alignment. Inspired by Social Choice Theory (Arrow, 2012), one future direction involves aggregating individual values into collective value agreements or judgments through democratic processes such as voting, discussion, and civic engagement (Gabriel, 2020). Therefore, building on the summary of human values in Section 4.1 and 5.2, future researchers can use democratic methods to identify a wide spectrum of human value subsets for AI alignment. Additionally, an important focus for future work lies in the curation of datasets that can represent these wide-spectrum human values, particularly in rich formats that go beyond simple ratings and rankings, which are commonly used in current alignment algorithms (e.g., RLHF (Ouyang et al., 2022), RLAIF (Bai et al., 2022b)). Another key research area to explore is around algorithms or methods (e.g., Bradley–Terry Model (Rafailov et al., 2023) or Elo Rating System (Bai et al., 2022b)), potentially drawing from economic theories, to convert heterogeneous human value data (e.g., principles or contextual information) into AI-compatible formats (e.g., scalars) for training reward models and guiding reinforcement learning. Also, future research can explore AI models that can act directly on aligning humans’ unstructured data, such as free-form descriptions of values, multimedia or sensor recordings depicting how people act in various situations, among others.
@@ 6.1.2. Aligning Humans to AI: Elicit nuanced and contextual human values during diverse interactions
Current alignment methods primarily capture human values using instructions, ratings and rankings, aiming to infer humans’ alignment goals (e.g., human values and intentions) in a data-driven manner. However, these limited data formats and interaction methods fall well short of fully specifying all of the relevant human values and constraints. Besides, when humans seek to collaborate with or utilize AI, they often encounter challenges in formulating optimal prompts, which require them to accurately and efficiently specify their requirements or expectations, in order to be able to use AI generated output or decision making in practice. Additionally, people often struggle to articulate all of their desired values accurately and comprehensively, especially as these values can dynamically change based on contextual and temporal situations. Relatedly, practices often overlook implicit human interactive signals that can also indicate values relevant to AI alignment.
Examples of Potential Future Solutions. One promising area for future research is optimizing interactive interfaces to elicit human values efficiently and effectively. These interfaces can leverage diverse interaction modes (see an initial summary in Figure 9) to fully capture human value information. For example, to identify a set of important principles for AI alignment, the interface can use the user’s conversational natural language specifications and generate a list of principle options for users to select from and/or revise. Recognizing that humans cannot always comprehensively articulate their values, another key research area is to develop interfaces that proactively interact with users. This could involve conversational techniques or scaffolding procedures to elicit nuanced, contextualized, and evolving information about an individual’s values. Additionally, to capture implicit human interaction signals that reflect human values for AI alignment, another area of future work involves designing systems that track interactions to develop real-time hypotheses about human values on these implicit signals, and then proactively validate them with humans as needed.
6.2. Challenge 2: Dynamic Co-evolution of Alignment
@@ 6.2.1. Aligning AI to Humans: Co-evolve AI with changes in humans and society
Existing literature often views AI alignment as a static process, neglecting the dynamic nature of alignment goals. However, a long-term perspective of alignment requires us to take into account the co-evolution of AI, humans, and society. On one hand, as AI systems scale up, they regularly acquire new and sometimes unexpected (”emergent”) capabilities, including learning from examples on the fly and adaptively pursuing goals. This leads to the challenge of ensuring that the goals AI might independently formulate remain consistently aligned with human values. Consequently, alignment solutions must involve continuous oversight and updates in response to AI’s evolution during development and deployment. On the other hand, AI advancements influence the state of the world, affecting human actions and even their underlying values and strategies (Hong et al., 2023). Thus, alignment solutions that continually adapt to the updated human and societal considerations may offer the most robust approach (Irving and Askell, 2019). Consequently, it is crucial to ensure that AI co-evolves with changes in human and societal values, cognition, and needs dynamically.
Examples of Potential Future Solutions. Updating AI to continuously adapt to the dynamic changes of humans and society requires adjustments in alignment data, models, and evaluation processes spanning both learning and deployment stages. Referring to the methodologies listed in Section 4.2, current strategies often rely on fine-tuning to update AI with a curated human value dataset during the learning process. However, this approach may lead to AI models forgetting or diminishing previous alignment goals. In situations where sufficient data for fine-tuning is unavailable, these methods becomes more challenging. Therefore, a crucial area for future research is to develop approaches for continually updating AI systems using limited data without compromising existing alignment values and performance. In real-world scenarios, it is common for AI systems to already be deployed at scale, making fine-tuning a challenging or impractical practice. Therefore, another important research direction involves adapting AI, during the deployment stage, without relying on fine-tuning. This challenge could potentially be addressed by forecasting the potential evolution trajectories of human values or behavioral patterns, and preparing AI with the flexibility to adapt in advance, for example, through prompting or intervention strategies. In summary, the co-evolution of AI with dynamic human and societal changes necessitates a synergetic collaboration among experts from various domains, such as machine learning and social sciences.
@@ 6.2.2. Aligning Humans to AI: Adapt humans and society to the latest AI advancements
As AI progresses, humans will likely need to transition from interacting with less capable systems to engaging advanced AI systems (e.g., AGI) (Morris et al., 2024). It’s imperative to develop interactive strategies which enable humans to effectively utilize AI across varying levels of model capability. While current AI systems still lag behind human performance in many tasks, questions persist about human resilience to AI mistakes. Efficiently identifying these mistakes and developing appropriate fallback actions, such as interactive intervention, remains crucial in various situations; identifying and handling errors, including knowing when to seek human intervention, will remain an important even for highly capable future systems. Looking ahead, as AI advances further, it becomes essential to develop interactive systems which enable humans to utilize AI that has capabilities surpassing their own. Research is needed to understand how individuals can interpret and validate the outputs of AI performing tasks beyond their current abilities, and also, how humans could leverage advanced AI in a sustainable way, i.e., benefiting their life without harmful side effects such as job displacement or loss of purpose. In addition, as AI becomes increasingly integrated into daily tasks, its influence on human values, behaviors, capabilities, and society remains uncertain. Thus, it’s vital to continuously examine the impact of AI advancements on individuals, social relationships, and broader societal changes.
Examples of Potential Future Solutions. While prior AI assistants have primarily aimed to complement human capabilities in task performance, future research on advanced AI systems should focus on developing validation mechanisms that enable humans to interpret and verify AI outputs. Inspired by initial explorations in Section 4.4.1, this could involve designing interactive interfaces that allow humans to request step-by-step justifications from AI, or integrating external tools to verify the truthfulness of AI outputs. This could also include developing interfaces that enable groups of humans to work together to validate AI outputs, since groups working jointly typically have broader and deeper ”intelligence” in practice than any single actor). Further, this approach can be extended to scalable validation tools, automating the verification process for large-scale applications such as in education and the workplace. Another important research direction is designing interactive strategies to enhance human capabilities by learning from advanced AI. This includes gaining knowledge and building technical and social skills. Furthermore, it is crucial to assess how humans and society adapt to AI advancements which, in turn, can guide the future evolution of AI. Potential research areas include evaluating changes in individual behavior and social relationships (e.g., how relationships with advanced AI may supplant human relationships), and societal governance (e.g., how legal and education systems may change as AI is used in place of traditional human skills). Examining these dynamic changes is essential for understanding the broader impact of AI advancements on humanity and society.
6.3. Challenge 3: Safeguarding Co-adaptation
As AI gains autonomy and capability, the risks associated with its instrumental actions, as a means toward accomplishing its final goals, increase. These actions can be undesirable for humans, i.e., power-seeking or evading human intervention. Therefore, safeguarding the co-adaptation between humans and AI is crucial. This can be achieved by empowering humans to understand and control AI’s instrumental goals, in order to ensure rational and ethical behaviors. We next explore future research to address this challenge from both directions.

### v3

@@ 5. Findings and Discussions on Current Gaps
In this section, our aim is to consolidate key findings derived from our analysis of the framework and the current state of literature we reviewed. We begin by analyzing the overall trends and gaps in the literature (Section 5.1). We then focus on three essential aspects: the relationship between human values and alignment (Section 5.2), the potential interaction modes used to specify human values (Section 5.3).
5.1. Meta Analysis of Trends and Gaps
Based on our coding of all papers, we computed the number of relevant papers for each dimension in the proposed bi-directional framework. Based on the distribution in Figure 4, we noticed that in existing literature certain dimensions are over- or under-represented. We outline the under-explored dimensions in two directions in the highlighted grey bars in the end of Section 4.2 and Section 4.4, respectively, and provide more details below.
Underexplored dimensions in AI-centered alignment research. Most literature specified human values using explicit human feedback, whereas implicit and simulated human feedback were largely under-explored. In the integrating human specifications to AI research question, Besides, developing and customizing AI models during the inference stage or using the interactive approach is under-explored. Also, human-in-the-loop evaluation is much less explored than automatic evaluation.
@@ 5.1. Meta Analysis of Trends and Gaps
Underexplored dimensions in Human-centered alignment research. Additionally, many studies on human cognitive adjustment to AI focus on enabling AI sensemaking and explanations so that humans can better understand, trust, and rely on AI. However, these studies often focused on explaining AI decision-making justification rather than educating people to acquire general skills and competencies required to understand, use, critique, and interact effectively with AI systems (i.e., AI literacy). AI literacy (Long and Magerko, 2020) plays a fundamental role in ensuring people understand and use AI correctly, and warrants deeper exploration. Also, we observed a wide range of studies (Wu et al., 2022; Gebreegziabher et al., 2023; Park et al., 2023; Ma et al., 2024) which developed interactive mechanisms and prototypes to empower human collaboration with AI. However, most studies assumed that AI plays an assistant role, being less capable than humans. This situation might change over time. Moreover, the influence of AI advancements on human behavior, social relationships, and societal changes is essential but remains largely unexplored.
| Takeaways of Interaction Techniques for Alignment. 1. Some common human feedback formats (rating, ranking) used in NLP/ML are not often studied in HCI. 2. Diverse human interactive feedbacks in HCI are not fully used in AI development in NLP/ML fields. |
5.2. Insights into Human Values for Alignment
@@
Our analysis, based on the adaptation of Schwartz’s Theory of Basic Values and our comprehensive literature review, identifies three critical findings for future research:
@@
Universal vs. Personalized AI Values. While certain values are universally expected from AI (e.g., capability, equity, responsibility), others may be undesirable in specific contexts (Ngo et al., 2024) (e.g., seeking power). Simultaneously, AI models should be adaptable to diverse human value systems (Sorensen et al., 2024). Research is needed to develop methods for identifying appropriate value sets for specific individuals or groups, and for customizing AI to align with user values while maintaining ethical principles.
Disparities in Value Expectations and Evaluation. The fundamental differences between humans and AI necessitate distinct approaches to value evaluation. For instance, assessing AI honesty may require mechanistic interpretability (Bereska and Gavves, 2024), a more rigorous standard than that applied to humans. Future studies should explore methods for evaluating and explaining AI values and calibrating human expectations accordingly.
5.3. Interaction Techniques for Specifying Human Values: A Cross-Domain Analysis
Our research reveals disparities in interaction techniques for human-AI value alignment across AI-centered (NLP/ML) and human-centered (HCI) domains. As depicted in Figure 5, this analysis focuses on three key areas:
Domain-Specific Interaction Techniques. The interaction techniques in AI-centered (NLP/ML) and Human-centered (HCI) alignment studies are often differ (Brath, 2021). NLP/ML studies primarily utilize numeric and natural language-based techniques. Also, NLP/ML research explore implicit feedback to extract human hidden feedback. In contrast, HCI research encompasses a broader range of graphical and multi-modal interaction signals (e.g., sketches, location information) beyond text and images. This disparity suggests potential gaps in extracting comprehensive human behavioral information.
Stage-Specific Interaction Techniques. In NLP/ML, the learning stage predominantly employs rating and ranking interactions for alignment in dataset generation. However, when humans use AI in the inference stage, as demonstrated in HCI research, involves more diverse user interactions. This discrepancy highlights the need for alignment between model development and practical deployment.
Divergent Data Utilization. NLP/ML typically uses interaction outputs as training datasets, while HCI analyzes this data to understand human behavior and feedback. As AI systems evolve, developing new interaction modes to capture a broader spectrum of human expression becomes crucial.
@@ 6. Future Directions
6.1. Specification Game
@@ 6.1.1. Aligning AI to Humans: Integrate fully specified human values into aligning AI
Individuals often possess value systems that encompass multiple values with varying priorities, rather than a single value, to guide their behaviors (Schwartz, 1994, 2012). Also, these priorities can change dynamically throughout an individual’s life stages. As such, It is more realistic to select values compatible with specific societies or situations, given the fact that we live in a diverse world (Gabriel, 2020). Future research, inspired by Social Choice Theory (Arrow, 2012), could focus on using democratic processes to aggregate individual values into collective agreements. Building on the summaries in Sections 4.1 and 5.2, researchers can employ democratic methods to identify diverse subsets of human values for AI alignment. Additionally, creating datasets that represent these values is crucial. Besides, it is crucial yet challenging for AI designers to investigate how to fully specify the appropriate values and to further integrate these values into AI alignment. Future important area involves developing algorithms, such as the Bradley-Terry Model (Rafailov et al., 2023) or Elo Rating System (Bai et al., 2022b), to convert heterogeneous human values into AI-compatible formats for training reward models and guiding reinforcement learning. Researchers should also explore AI models capable of aligning with unstructured human data, including free-form descriptions of values, multimedia, or sensor recordings depicting human behavior.
@@ 6.1.2. Aligning Humans to AI: Elicit nuanced and contextual human values during diverse interactions
Current alignment methods use instructions, ratings, and rankings to infer human values, which can not fully capture all relevant human values and constraints. Future research should focus on optimizing interactive interfaces to efficiently elicit human values. These interfaces can leverage diverse interaction modes to capture comprehensive human value information. Additionally, people often struggle to formulate optimal prompts for AI, accurately specify their requirements, and articulate their desired values, which can change based on context and time. Developing proactive interfaces that use conversational techniques to elicit nuanced and evolving values is also crucial. Implicit human signals that indicate values are also frequently overlooked. Additionally, systems that track interactions to hypothesize and validate implicit human values in real-time should be designed.
6.2. Dynamic Co-evolution of Alignment
@@ 6.2.1. Aligning AI to Humans: Co-evolve AI with changes in humans and society
(i) Existing literature often treats AI alignment as static, ignoring its dynamic nature. A long-term perspective must consider the co-evolution of AI, humans, and society. As AI systems evolve and scale up, they gain new capabilities, making it essential to ensure their goals remain aligned with human values. Thus, alignment solutions require continuous oversight and updates. Future research should develop methods for continuously updating AI with limited data without compromising alignment values and performance. This could involve forecasting human value evolution and preparing AI with flexible strategies like prompting or interventions. (ii) Additionally, AI advancements also influence human actions and values, necessitating adaptive alignment solutions. Ensuring AI co-evolves with human and societal changes is crucial for robust alignment. This challenge could potentially be addressed by forecasting the potential evolution trajectories of human values or behavioral patterns, and preparing AI with the flexibility to adapt in advance, for example, through prompting or intervention strategies.
@@ 6.2.2. Aligning Humans to AI: Adapt humans and society to the latest AI advancements
(i) While current AI systems lag behind humans in many tasks, identifying and handling AI mistakes, including knowing when to seek human intervention, remains essential. Future research should focus on developing validation mechanisms that enable humans to interpret and verify AI outputs. This could involve designing interfaces that allow humans to request step-by-step justifications from AI or integrating tools to verify the truthfulness of AI referring to Section 4.4.1. Additionally, developing interfaces that enable groups of humans to collaboratively validate AI outputs and creating scalable validation tools for large-scale applications are important directions. (ii) As AI advances, it becomes essential to develop systems that enable humans to utilize AI with capabilities surpassing their own. Research is needed to understand how individuals can interpret and validate AI outputs for tasks beyond their abilities and leverage advanced AI sustainably, avoiding issues like job displacement or loss of purpose. Another research direction is designing strategies to enhance human capabilities by learning from advanced AI, including gaining knowledge and building skills. (iii) As AI integrates more into daily tasks, its impact on human values, behaviors, capabilities, and society remains uncertain. Continuous examination of AI’s influence on individuals, social relationships, and broader societal changes is vital. Research should assess how humans and society adapt to AI advancements, guiding AI’s future evolution. Potential areas include evaluating changes in individual behavior, social relationships, and societal governance as AI replaces traditional human skills. Understanding these dynamic changes is essential for grasping the broader impact of AI on humanity and society.
6.3. Safeguarding Co-adaptation
As AI gains autonomy and capability, the risks associated with its instrumental actions, as a means toward accomplishing its final goals, increase. These actions can be undesirable for humans. Therefore, safeguarding the co-adaptation between humans and AI is crucial. We next explore future research to address this challenge from both directions.